need help urgent!
write the linear function whose graph has the given slope (m) and passes through the given point.
1. m = 1; (-2,-3)
2. m = -4;(0,-3)
write the linear function whose graph passes through the points.
3. (-3,6)(-5,9)
4. (-5,3)(0,6)
write the linear function whose slope y intercept are given.
5. slope 1; y intercept -8
6. slope 4; y intercept -4
write the linear function whose x intercept a and y intercept b are given.
7. a=1; b=5
8. a=3; b=1/2
9-10 Mr. Reyman used 3.2 m of copper wire which weighs 0.45 kg, How much will 6m of copper wire weigh?

Answer by Alan3354(69443)

write the linear function whose graph has the given slope (m) and passes through the given point.
1. m = 1; (-2,-3)
Use y-y1 = m*(x-x1) where (x1,y1) is the point
y+3 = 1*(x+2)
y+3 = x+2
x-y = 1
-------------
2. m = -4;(0,-3)
write the linear function whose graph passes through the points.
3. (-3,6)(-5,9)
Find the slope
m = diffy/diffx = (6-9)/(-3+5)
m = -3/2
Do the same as #1, using either point.
----------------------------------------
4. (-5,3)(0,6)
write the linear function whose slope y intercept are given.
5. slope 1; y intercept -8
y = mx + b (b is the y-intercept)
y = x-8
-----------
6. slope 4; y intercept -4
write the linear function whose x intercept a and y intercept b are given.
7. a=1; b=5
m = -b/a = -5
Then it's the same as #1, using either point, (1,0) or (0,5).
8. a=3; b=1/2
9-10 Mr. Reyman used 3.2 m of copper wire which weighs 0.45 kg, How much will 6m of copper wire weigh?
It weighs 0.45/3.2 kg/m.
6*0.45/3.2 = 0.84375 kg